Assign an oxidation state to each atom in each species.

Part A

Br2

Express your answer as an integer with the appropriate sign.

Part B

Cr3+

Express your answer as an integer with the appropriate sign.

Part C

CF4

Express your answers as integers with the appropriate signs separated by a comma.

C, F:

Part D

ZnCl2

Express your answers as integers with the appropriate signs separated by a comma.

Zn, Cl:

Part E

P2O3

Express your answers as integers with the appropriate signs separated by a comma.

P, O:

Part F

NO3−

Express your answers as integers with the appropriate signs separated by a comma.

N, O:
Answer #1

Part A

Br2 ( Its a molecule, the oxidation number of an atom is zero in a neutral substance that contains atoms of only one element)
Hence Oxidation state of Br2 is Zero; "0"

Part B

Cr3+ ( Here Oxidation state of Cr is +3)

Part C

CF4 ( Here Oxidation state of C is +4 , Oxidation State of F is -1 )

Part D

ZnCl2 ( Here Oxidation state of Zn is +2 , Oxidation state of Cl is -1 )

Part E

P2O3 ( here oxidation state of P is +3 ; Oxidation state of O is -2)

Part F

NO3 ( here oxidation state of N is +5 ; Oxidation state of O is -2 )
X-3(-2) = -1
X= -1 +6 = = +5 ( where X is oxidation of N )

You have to remember always , Oxide O oxidation state is -2 , Halogens Oxidation state is -1